# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ C.4.11 Proof of Lemma 25
$\bigcap_{S\in[p],|S|\leq\sqrt{p}}\bigcap_{1\leq k\neq l\leq K}\mathcal{C}_{N,kl}^{S}$
@@
$\mathbb{P}\bigl{(}\bigcup_{S\in[p],|S|\leq\sqrt{p}}\bigcup_{1\leq k\neq l\leq K}(\mathcal{C}_{N,kl}^{S})^{C}\bigr{)}$ $\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{N,kl}^{S})^{C})$
@@
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}G^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}$
@@ C.4.12 Proof of Lemma 26
Since the sum of inner products of Gaussian random vectors is challenging to analyze directly, we first condition on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$ and then apply the concentration inequality for the supremum of a Gaussian process (Lemma 9). Fix $(k,l)$ such that $1\leq k\neq l\leq K$ and $S\subset[p]$ such that $|S|\leq\sqrt{p}$. Conditioned on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$, $G^{S}_{2,kl}(\mathbf{v})$ can be viewed as a canonical Gaussian process $\langle G,\mathbf{w}\rangle$, where $G\sim\mathcal{N}(0,\mathbf{I}_{|G_{k}^{\ast}|})$ and $\mathbf{w}\in\mathbb{W}_{k}$, with
@@
[…]
@@ 2. Chi-square tail bound
Now we use the Chi-square tail inequality to further bound ¯$\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}$ and ¯$\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}$. If we define the following event:
@@
[…]
@@ Conclusion
=$\overset{(i)}{=}\sum_{1\leq k\neq l\leq K}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{N^{S}_{kl}(\mathbf{v},\mathbf{X})\;N^{S}_{lk}(\mathbf{v},\mathbf{X})}{t^{S}_{lk}(\dot{\lambda}^{S})}$
@@
$\overset{(i)}{\lesssim}\sum_{1\leq k\neq l\leq K}\biggl{\{}\frac{\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}}}}$
×$\hskip 60.00009pt\times\sqrt{|G_{k}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{l}^{\ast}|+\sqrt{\xi^{S}}}\biggr{\}}$
@@
Now, for the second term, we first recall from Theroem 1 that
$m=2\min_{1\leq k\neq l\leq K}\left\{\frac{|G_{k}^{\ast}||G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\}.$
Then we have:
@@ Appendix D Proof of Theorem2
~$\tilde{\mathcal{M}}(s,\rho,2):=\bigl{\{}(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}):$ $\boldsymbol{\mu}_{1}=\boldsymbol{\mu},\boldsymbol{\mu}_{2}=-\boldsymbol{\mu},\boldsymbol{\mu}\in\mathbb{R}^{p},\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}^{2}\geq\rho^{2},$
@@
=$~{}\overset{(ii)}{=}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}1-(1-\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}})^{n}$
=$~{}\overset{(iii)}{=}1-\bigl{(}1-\Phi(-\rho/\sigma)\bigr{)}^{n},$ (67)
@@ E.1 Derivation of the SDP objective function
This section presents the derivation of the SDP objective (1), which was omitted from Section 2.1 for brevity. The derivation follows Sections 2 and 3 of Zhuang et al., (2023) and Section 1 of Chen and Yang, (2021), and is included here for completeness. We proceed in the following order:
@@
If $\boldsymbol{\Sigma}^{\ast}=\mathbf{I}_{p}$, the maximization of (70) is formulated as the following K-means problem, by removing the constant factor $-1/2$:
@@ E.2 Intuition for the SDP objective in Algorithm 1
Appendix F Details for Iterative Algorithms
This section supplements the iterative algorithms presented in Section 3 by providing omitted details. Appendix F.1 derives the uniform bound presented in (8), which is used in the variable selection step of Algorithm 2. Appendix F.2 provides additional details for the ISEE subroutine, which plays a central role in Algorithm 3.
F.1 Derivation of the uniform bound in the variable selection step of Algorithm 2
@@ F.2.1 Regression relationship on the data sub-vectors
[…]
@@
$-\boldsymbol{\Sigma}_{A,A^{c}}^{\ast}(\boldsymbol{\Sigma}_{A^{c},A^{c}}^{\ast})^{-1}\overset{(i)}{=}-\mathbf{B}\mathbf{D}^{-1}\overset{(ii)}{=}\mathbf{W}^{-1}(-\mathbf{W}\mathbf{B}\mathbf{D}^{-1})\overset{(iii)}{=}(\boldsymbol{\Omega}_{A,A}^{\ast})^{-1}\boldsymbol{\Omega}_{A,A^{c}}^{\ast},$
@@ F.2.2 Key identities on transformed cluster center and noise
(~$(\tilde{\mathcal{E}}_{i})_{A}=\boldsymbol{\Omega}_{A,A}\mathbf{E}_{A,i},$
@@
~$\tilde{\mathcal{E}}_{A,i}$ $=\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\mathbf{X}_{i}-\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\boldsymbol{\mu}^{\ast}_{1}$
@@ F.2.3 Estimation error bound for the ISEE subroutine
=$\;\overset{(i)}{=}\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$
=$\;\overset{(ii)}{=}\bigl{\|}\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}~{}+~{}\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\bigr{\|}_{2}$
@@
Under some regularity conditions, Fan and Lv, (2016) proves that $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$, $\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$, and $\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}-(\boldsymbol{\alpha}_{2})_{A_{\ell}}\|_{2}$ all have have convergence rates of order
$\max\{J\log p/n,\sqrt{\log p/n}\}.$
Therefore, we can conclude that
@@
8:     $\hat{\mathbf{E}}_{A_{\ell},\cdot}\leftarrow(\hat{\mathbf{E}}_{A_{\ell},1},\ldots,\hat{\mathbf{E}}_{A_{\ell},n})$triangleright Residual block
9:     $\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}\leftarrow(\frac{1}{n}\hat{\mathbf{E}}_{A_{\ell},\cdot}\hat{\mathbf{E}}_{A_{\ell},\cdot}^{\top})^{-1},$triangleright Precision matrix block
10:     $(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}},$triangleright Transformed cluster 1 center block
11:     $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}$triangleright Transformed cluster 2 center block
@@
13:     $\hat{\tilde{\mathbf{X}}}_{A_{\ell},\cdot}\leftarrow\hat{\tilde{\mathbf{M}}}_{A_{\ell},\cdot}+\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}\hat{\mathbf{E}}_{A_{\ell},\cdot}$triangleright Transformed data block
14:$\hat{\tilde{\boldsymbol{\mu}}}_{1}\leftarrow\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{1}}^{\top},\ldots,(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{L}}^{\top}\bigr{)}^{\top},$triangleright Aggregate cluster 1 center blocks
15:$\hat{\tilde{\boldsymbol{\mu}}}_{2}\leftarrow\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{1}}^{\top},\ldots,(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{L}}^{\top}\bigr{)}^{\top}$triangleright Aggregate cluster 2 center blocks
@@
Appendix G Supplementary details for numerical study

### v2

@@ C.4.11 Proof of Lemma 25
$\bigcap_{S\in[p],|S|\leq\sqrt{p}}\bigcap_{1\leq k\neq l\leq K}\hskip-10.00002pt\mathcal{C}_{N,kl}^{S}$
@@
$\mathbb{P}\bigl(\hskip-10.00002pt\bigcup_{S\in[p],|S|\leq\sqrt{p}}\bigcup_{1\leq k\neq l\leq K}\hskip-10.00002pt(\mathcal{C}_{N,kl}^{S})^{C}\bigr)$ $\leq\hskip-3.99994pt\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{N,kl}^{S})^{C})$
@@
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-10.00002ptG^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}$
@@ C.4.12 Proof of Lemma 26
Since the sum of inner products of Gaussian random vectors is challenging to analyze directly, we first condition on $\overline{\mathcal{E}}_{S,l}^{\ast}$ and then apply the concentration inequality for the supremum of a Gaussian process (Lemma 9). Fix $(k,l)$ such that $1\leq k\neq l\leq K$ and $S\subset[p]$ such that $|S|\leq\sqrt{p}$. Conditioned on $\overline{\mathcal{E}}_{S,l}^{\ast}$, $G^{S}_{2,kl}(\mathbf{v})$ can be viewed as a canonical Gaussian process $\langle G,\mathbf{w}\rangle$, where $G\sim\mathcal{N}(0,\mathbf{I}_{|G_{k}^{\ast}|})$ and $\mathbf{w}\in\mathbb{W}_{k}$, with
@@
$\mathrm{EC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})\lesssim\int_{0}^{\infty}\sqrt{\log N(\mathbb{W}_{k},\epsilon)}\,d\epsilon$ $\overset{(i)}{\lesssim}\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}\int_{0}^{\infty}\sqrt{|G_{k}^{\ast}|\log(1/\varepsilon)}\,d\varepsilon$ $=\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}\sqrt{|G_{k}^{\ast}|},$
@@ 2. Chi-square tail bound
Now we use the Chi-square tail inequality to further bound $\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}$ and $\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}$. If we define the following event:
@@
$\overset{(i)}{\leq}\left(\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{E}\bigl[\mathrm{CP}_{N,kl}^{S}\;\mathbbm{1}(\mathcal{B}^{S}_{N,kl})\bigr]\right)+\frac{K^{2}\pi^{2}}{6n}$
@@ Conclusion
$\overset{(i)}{=}\sum_{1\leq k\neq l\leq K}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{N^{S}_{kl}(\mathbf{v},\mathbf{X})\;N^{S}_{lk}(\mathbf{v},\mathbf{X})}{t^{S}_{lk}(\dot{\lambda}^{S})}$
@@
$\overset{(i)}{\lesssim}\sum_{1\leq k\neq l\leq K}\frac{\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}}}\sqrt{|G_{k}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{l}^{\ast}|+\sqrt{\xi^{S}}}}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl(\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr)}$
@@
Now, for the second term, we first recall from Theroem 1 that $m=2\min_{1\leq k\neq l\leq K}\left\{\frac{|G_{k}^{\ast}||G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\}$. Then we have:
@@ Appendix D PROOF OF THEOREM 2
$\tilde{\mathcal{M}}(s,\rho,2):=\bigl\{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}):$ $\boldsymbol{\mu}_{1}=\boldsymbol{\mu},\boldsymbol{\mu}_{2}=-\boldsymbol{\mu},\boldsymbol{\mu}\in\mathbb{R}^{p},\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}^{2}\geq\rho^{2},$
@@
$~\overset{(ii)}{=}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}1-(1-\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}})^{n}$
$~\overset{(iii)}{=}1-\bigl(1-\Phi(-\rho/\sigma)\bigr)^{n},$ (71)
@@ E.1 Derivation of the SDP objective function
This section presents the derivation of the SDP objective (3), which was omitted from Section 2 for brevity. The derivation follows Sections 2 and 3 of Zhuang et al., (2023) and Section 1 of Chen and Yang, (2021), and is included here for completeness. We proceed in the following order:
@@
If $\boldsymbol{\Sigma}^{\ast}=\mathbf{I}_{p}$, the maximization of (74) is formulated as the following $K$-means problem, by removing the constant factor 1/2-1/2:
@@ E.2 Intuition for the SDP objective in Algorithm 1
Appendix F DETAILS FOR PROPOSED ITERATIVE ALGORITHMS
This section supplements the iterative algorithms presented in Section 3 by providing omitted details. Appendix F.1 derives the uniform bound presented in (11), which is used in the feature selection step of Algorithm 2. Appendix F.2 provides additional details for the ISEE subroutine, which plays a central role in Algorithm 3.
F.1 Derivation of the uniform bound in the feature selection step of Algorithm 2
@@ F.2.1 Regression relationship on the data sub-vectors
$\hskip-16.99998pt\underbrace{\mathbf{X}_{A,i}}_{\hskip 9.01396pt\text{response}\in\mathbb{R}^{|A|}}\hskip-6.99997pt=\hskip 6.99997pt\underbrace{(\boldsymbol{\mu}_{1}^{\ast})_{A}+(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}}_{:=(\boldsymbol{\alpha}_{1})_{A}\in\mathbb{R}^{|A|}~\text{(intercept)}}~-~\underbrace{(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}}_{\hskip-21.30568pt\text{slope}\in\mathbb{R}^{|A|\times(p-|A|)}}\hskip-16.00008pt\underbrace{\mathbf{X}_{A^{c},i}}_{\hskip 9.01396pt\text{predictor}\in\mathbb{R}^{p-|A|}}\hskip-13.00005pt+\hskip-15.3pt\underbrace{\mathbf{E}_{A,i}}_{\hskip 17.20844pt\text{residual}\in\mathbb{R}^{|A|}}\hskip-21.00009pt,$
@@
$\boldsymbol{\Sigma}_{A,A^{c}}^{\ast}(\boldsymbol{\Sigma}_{A^{c},A^{c}}^{\ast})^{-1}\overset{(i)}{=}-\mathbf{B}\mathbf{D}^{-1}\overset{(ii)}{=}\mathbf{W}^{-1}(-\mathbf{W}\mathbf{B}\mathbf{D}^{-1})\overset{(iii)}{=}(\boldsymbol{\Omega}_{A,A}^{\ast})^{-1}\boldsymbol{\Omega}_{A,A^{c}}^{\ast},$
@@ F.2.2 Key identities on transformed cluster center and noise
$(\tilde{\mathcal{E}}_{i})_{A}=\boldsymbol{\Omega}_{A,A}\mathbf{E}_{A,i},$
@@
$\tilde{\mathcal{E}}_{A,i}$ $=\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\mathbf{X}_{i}-\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\boldsymbol{\mu}^{\ast}_{1}$
@@ F.2.3 Estimation error bound for the ISEE subroutine
$\;\overset{(i)}{=}\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$
$\;\overset{(ii)}{=}\bigl\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}~+~\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\bigr\|_{2}$
@@
Under some regularity conditions, Fan and Lv, (2016) proves that $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$, $\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$, and $\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}-(\boldsymbol{\alpha}_{2})_{A_{\ell}}\|_{2}$ all have have convergence rates of order $\max\{J\log p/n,\sqrt{\log p/n}\}$. Therefore, we can conclude that
@@
8:  $\hat{\mathbf{E}}_{A_{\ell},\cdot}\leftarrow(\hat{\mathbf{E}}_{A_{\ell},1},\ldots,\hat{\mathbf{E}}_{A_{\ell},n})$ $\triangleright$ Residual block
9:  $\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}\leftarrow(\frac{1}{n}\hat{\mathbf{E}}_{A_{\ell},\cdot}\hat{\mathbf{E}}_{A_{\ell},\cdot}^{\top})^{-1},$ $\triangleright$ Precision matrix block
10:  $(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}},$ $\triangleright$ Transformed cluster 1 center block
11:  $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}$ $\triangleright$ Transformed cluster 2 center block
@@
13:  $\hat{\tilde{\mathbf{X}}}_{A_{\ell},\cdot}\leftarrow\hat{\tilde{\mathbf{M}}}_{A_{\ell},\cdot}+\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}\hat{\mathbf{E}}_{A_{\ell},\cdot}$ $\triangleright$ Transformed data block
[…]
@@
Appendix G DETAILS FOR NUMERICAL STUDY